OpenStudy (anonymous):

I've got two equations but i cant replace anything in one for the other, how do i figure this out? A carpenter is building a rectangular room with a fixed perimeter of 356 ft. What dimensions would yield the max. area? What is the max area?

OpenStudy (anonymous):

i think the maximum area is a square

OpenStudy (anonymous):

There are several ways of doing this, what course is this from?

OpenStudy (anonymous):

Algebra 2

OpenStudy (radar):

Assuming jimmyrep is correct then the sides would be 89 (356/4)

OpenStudy (anonymous):

then my area would be 7921?

OpenStudy (radar):

Yes (assuming jimmyrep is correct)

OpenStudy (anonymous):

okay thank you. im just terrible with story problems they catch me off guard

OpenStudy (radar):

I would wait and see what Nancy Lam has to say

OpenStudy (anonymous):

will do

OpenStudy (anonymous):

P= (L+W)2 365= (L+W)2 365/2= L+W 182.5 = L+W 182.5/2 =91.25 \[91.25^{2}=8.327?\]

OpenStudy (radar):

Nancy don't you mean 356?

OpenStudy (anonymous):

ok i was wron number, any way use 356 instate 365 my equation is right?

OpenStudy (radar):

I was with you until the 5th line. what is that about???

OpenStudy (radar):

Oh, I see you are letting L=W so it would be square. Seems like a burdensome method when you coluld just divide 356 by 4

OpenStudy (anonymous):

P= (L+W)2 356= (L+W)2 356/2= L+W 178 = L+W 178/2 =89 \[89^{2}=7.921\]

OpenStudy (radar):

It appears that Nancy Lam agrees with the 89 answer.

OpenStudy (anonymous):

check: P= (L+W)2 356=(89+89)2 356=178*2 356=356

OpenStudy (anonymous):

Thank

OpenStudy (radar):

You realize niki 150 that this is true, if and only if, the maximum area is a square. I hope jimmyrep is right.

OpenStudy (anonymous):

yeah i know. but there is no other way to figure it out so it's gotta be a square. if there was another way there isnt enough information to do so.

OpenStudy (radar):

Agree, good luck with your studies. Have a good weekend

OpenStudy (anonymous):

thanks.

OpenStudy (anonymous):

u r we com